According to a government study among adults in the 25- to 34-year age group, the mean amount spent per year on reading and entertainment is $2,125. Assume that the distribution of the amounts spent follows the normal distribution with a standard deviation of $439. (Round your z-score computation to 2 decimal places and final answers to 2 decimal places.)

What percent of the adults spend more than $2,400 per year on reading and entertainment?

What percent spend between $2,400 and $3,400 per year on reading and entertainment?

What percent spend less than $1,000 per year on reading and entertainment?
Answer #1

439 1 Ginen M= 2125 f r = 439. (6) p (272400) => p (9s 2400 -2105) plaz0.6264) =0.5 - P (OKU 20.6264) = 0.5 – 0.2357 :p (2x24(c) plaal000) = P(R=< 100 101 125) Plze-2-5626) = 10.5 - P(-2.50 2220) 2 0.5-0.4948 - 2,500 llence Plax1000) = 0.0052 0.524.